# Approximate helical symmetry in compact binaries

### v1

@@ Abstract
The inspiral of a circular, non-precessing binary exhibits an approximate helical symmetry. The effects of eccentricity, precession, and radiation reaction break the exact symmetry. We estimate the failure of this symmetry using the flux of the BMS charge corresponding to helical symmetry carried away by gravitational waves. We analytically compute the helical flux for binaries moving on eccentric orbits and quasi-circular orbits without precession using post-Newtonian theory. The helical flux is non-vanishing at the 0PN order for eccentric orbits as expected. We analytically predict the helical flux to be at a relative 5PN order for quasi-circular non-precessing binaries. This prediction is compared with 113 quasi-circular non-precessing numerical relativity waveforms from the SXS catalog. We find good agreement between analytical and numerical results for quasi-circular non-precessing binaries establishing that helical symmetry starts to break at 5PN for these binaries.
@@ II Helical symmetry
For asymptotically flat spacetimes, the asymptotic symmetry group at $\mathscr{I}^{+}$ is not the Poincaré group, but an enlarged group called the Bondi–van der Burg–Metzner–Sachs (BMS) group [14, 15, 16]. If $\Omega$ were simply constant, then $K^{\alpha}\partial_{\alpha}$ would be a BMS generator. The flux at $\mathscr{I}^{+}$ due to a BMS vector field $\xi$ is [17]
$F_{\xi}=-\frac{1}{32\pi}\int d^{2}x\,N^{ij}\mathcal{L}_{\xi}C_{ij},$ (2)
where $C_{ij}$ is the Bondi news aspect and $N_{ij}$ is the Bondi news. The helical vector field can be expressed in Bondi coordinates as
@@ III.1 Eccentricity
$=\frac{352c^{5}\nu^{2}x^{5}e^{2}}{15G}+\mathcal{O}\left(e^{2}\right).$ (16b)
@@ III.2.1 Post-Newtonian helical flux
The helical flux requires the computation of energy and angular momentum flux as shown in Eq. (5). The energy flux has been computed in the literature up to 4.5PN order. This computation begins with constructing the metric perturbation in a Bondi-type coordinate system $(u,\mathbf{X})$,
@@
$-\frac{2\ell}{c(\ell+1)}N_{kL-2}\epsilon_{kl(i}\text{V}_{j)lL-2}(u)\bigg{]}+\mathcal{O}\left(\frac{1}{R^{2}}\right)\,$
where $N^{i}=X^{i}/R$ and $\perp_{abij}$ is the projection tensor. The metric perturbation is expressed in terms of radiative mass-type $\text{U}_{L}$ and current-type $\text{V}_{L}$ multipole moments in a transverse-traceless gauge. It is straightforward to obtain the multipole expansion of the energy flux using Eq. (2),
@@
We outline the procedure to compute the hereditary integrals involved in this computation in the appendix. Using these flux expressions, we compute the helical flux as expressed in Eq. (5).
@@ IV.0.3 Results
Figure 4: The instantaneous PN order computed for 113 SXS simulations. The curves in the foreground are for a quasi-circular non-precessing simulation, SXS:BBH:1154. The helical flux is computed using both $\Omega_{\text{rot}}$ (purple) and $\Omega_{(2,2)}$ (green). For reference, the yellow and blue curves correspond to the instantaneous PN orders of $\dot{E}$ and $\Omega\cdot\dot{L}$ respectively.
@@ V Discussion and conclusions
For quasi-circular non-precessing binary systems, we analytically predicted that the dominant contribution to the helical flux comes from the quadrupole moment. The PN order of the helical flux starts at a 5PN order relative to its contribution to the energy flux. The analytical prediction agrees with our numerical computation of the instantaneous PN order of the helical flux. A robust analytical comparison requires PN results that are not yet available in the literature. We leave the computation of this exact analytical expression of helical flux for future work.

### v2

@@ Abstract
The inspiral of a circular, non-precessing binary exhibits an approximate helical symmetry. The effects of eccentricity, precession, and radiation reaction break the exact symmetry. We estimate the failure of this symmetry using the flux of the Bondi–van der Burg–Metzner–Sachs (BMS) charge corresponding to helical symmetry carried away by gravitational waves. We analytically compute the helical flux for binaries moving on eccentric orbits and quasi-circular orbits without precession using post-Newtonian theory. The helical flux is non-vanishing at the 0PN order for eccentric orbits as expected. We analytically predict the helical flux to be at a relative 5PN order for quasi-circular non-precessing binaries. This prediction is compared with 113 quasi-circular non-precessing numerical relativity waveforms from the SXS catalog. We find good agreement between analytical and numerical results for quasi-circular non-precessing binaries establishing that helical symmetry starts to break at 5PN for these binaries.
@@ II Helical symmetry
For asymptotically flat spacetimes, the asymptotic symmetry group at $\mathscr{I}^{+}$ is not the Poincaré group, but an enlarged group called the Bondi–van der Burg–Metzner–Sachs (BMS) group [14, 15, 16]. The BMS group is a semidirect product between the Poincaré group and supertranslations, which are interpreted as direction-dependent translations. Supertranslations form an infinite-dimensional Abelian subgroup. Every generator in the BMS algebra, given by a vector field $\xi$, gives rise to an associated flux at $\mathscr{I}^{+}$ [17, 18]
$F_{\xi}=-\frac{1}{32\pi}\int d^{2}x\,N^{AB}\delta_{\xi}C_{AB},$ (2)
where $C_{AB}$ is the shear and $N_{AB}$ is the Bondi news. Here $\delta_{\xi}C_{AB}$ is the change in the shear resulting from $g_{ab}\to g_{ab}+\mathcal{L}_{\xi}g_{ab}$ [18].
When $\Omega$ is time-dependent, $K^{\alpha}\partial_{\alpha}$ is not in the BMS algebra. However, when $\Omega$ is constant, we have a helical symmetry and $K^{\alpha}\partial_{\alpha}$ is a BMS generator (specifically, a Poincaré generator). The helical vector field can be expressed in Bondi coordinates as
@@ III.1 Eccentricity
$=\frac{352c^{5}\nu^{2}x^{5}e^{2}}{15G}+\mathcal{O}\left(e^{4}\right).$ (16b)
@@ III.2.1 Post-Newtonian helical flux
The helical flux requires the computation of energy and angular momentum flux as shown in Eq. (5). The energy flux has been computed in the literature up to 4.5PN order. This computation begins with constructing the metric perturbation in a Bondi-type coordinate system $(u,\mathbf{X})$
@@
$-\frac{2\ell}{c(\ell+1)}N_{kL-2}\epsilon_{kl(i}\text{V}_{j)lL-2}(u)\bigg{]}+\mathcal{O}\left(\frac{1}{R^{2}}\right),\,$
where $N^{i}=X^{i}/R$ and $\perp_{abij}$ is the projection tensor. The metric perturbation is expressed in terms of radiative mass-type $\text{U}_{L}$ and current-type $\text{V}_{L}$ multipole moments in a transverse-traceless gauge. It is straightforward to obtain the multipole expansion of the energy flux using Eq. (2), and is expressed as
@@
We outline the procedure to compute the hereditary integrals involved in this computation in the appendix. Using these flux expressions, we compute the helical flux as expressed in Eq. (5),
@@ IV.0.3 Results
Figure 4: The instantaneous PN order computed for 113 SXS simulations. The curves in the foreground are for a quasi-circular non-precessing simulation, SXS:BBH:1154. The helical flux is computed using both $\Omega_{\text{rot}}$ (purple) and $\Omega_{(2,2)}$ (green). For reference, the yellow and blue curves correspond to the instantaneous PN orders of $\dot{E}$ and $\Omega\cdot\dot{L}$ respectively. The curvature near the edges is a numerical artifact of the moving average procedure.
@@ V Discussion and conclusions
For quasi-circular non-precessing binary systems, we analytically predicted that the dominant contribution to the helical flux comes from the quadrupole moment. The PN order of the helical flux starts at a 5PN order relative to its contribution to the energy flux. The analytical prediction agrees with our numerical computation of the instantaneous PN order of the helical flux. A robust analytical comparison requires PN results that are not yet available in the literature. We leave computing the exact analytical expression of helical flux to future work. Our numerical analysis justifies previous intuition that circular orbits remains circular up to 5PN order during the inspiral phase under radiation reaction [13]. Semi-analytic waveform models, such as EOB or IMRPhenom, should satisfy the approximate helical symmetry to a degree consistent with NR waveforms to ensure their accuracy. The helical flux can therefore act as a consistency check on these waveform models.